A company has the following capital structure to support its $600,000 in Total Assets. What percentage of the capital is Common Equity?

$
Debt 300,000
Common Stock 100,000
Retained Earnings 200,000
Answer #1

Equity = Common stock + retained earnings

Equity = 100,000 + 200,000 = 300,000

Percentage of common equity = (Common stock / total capital structure) * 100

Percentage of common equity = (300,000 / 600,000 ) * 100

Percentage of common equity = 50%